# Neural Exploratory Landscape Analysis

### v2

@@ A.1 Technical Details of $\Lambda_{\theta}$
Two-stage Attention Block. We construct a two-stage attention block (Ts-Attn) to aggregate optimization status information across candidate solutions and across each dimension of the decision variables. The overall computation graph of the Ts-Attn is illustrated in the right of Figure 2, of which a basic component is the attention block ($Attn$). As illustrated in the left of Figure 2, the $Attn$ block mainly follows designs of the original Transformer [25], except that the layer normalization [53] is used instead of batch normalization [54]. Given a group of $L$ input encoding vectors $X_{in}\in\mathbb{R}^{L\times h}$, Eq. (5) details the computation of the $Attn$ block.
@@
where MHSA, LN and FF denote the multi-head self-attention [25] (with the hidden dimension of $h$), layer normalization [53] and linear feed forward layer respectively. The output $o$ holds identical shape with the input $X_{in}$. In our Ts-Attn block, we employ an $Attn$ block $Attn_{inter}$ for the first cross-solution information sharing stage, and the other $Attn$ block $Attn_{intra}$ for the second cross-dimension information sharing stage (illustrated in the right of Figure 2). The Ts-Attn receives the input encoding $s$ of the current candidate population, and then advances the information sharing in both cross-solution and cross-dimension level. The computation is detailed in Eq. (6).
@@
At the first stage, we let the input encoding $s$ (attained from the pre-processing module) pass through $Attn_{inter}$. Since we group the encodings of the same dimension of all candidates in $s$, the $Attn_{inter}$ promotes the optimization information sharing across candidates in current population. From the first stage, we obtain a group of hidden features $H$ with the shape of $d\times m\times h$. At the second stage, we first transpose $H$ into the shape of $m\times d\times h$ to regroup all dimensions of a candidate together. We then add $cos$/$sin$ positional encoding (PE) over the transposed $H$ to inform the order of different dimensions in a candidate. We then let $H$ pass through $Attn_{intra}$ to advance the information sharing among the different dimensions within the same candidate. The output of $Attn_{intra}$ holds the shape of $m\times d\times h$. At last, we apply MeanPooling on $H_{out}$ to get the landscape feature for each candidate $F_{indiv}$ in the population, and apply a second MeanPooling on $F_{indiv}$ to get the landscape feature for the whole candidate population $F_{pop}$. We have to note that we calculate both $F_{indiv}$ and $F_{pop}$ to make our NeurELA compatible with diverse MetaBBO algorithms, which either require the landscape feature of the whole population (e.g., [23]) or require a separate landscape feature for each candidate (e.g., [21]). The highly parallelizable attention-based neural-network architecture ensure the scalability of our method as the amount of the sampled candidates or the dimensions of the problem increases.
@@ A.2 Train-test split of BBOB testsuites
BBOB contains 24 synthetic problems owning various landscape properties [15] (e.g. multi-modality, global structure, separability and etc.). Table 2 list all of the 24 problems according to their category. Due to the diversity of properties, how to split these problems into train-test set becomes a key issue to ensure the training performance and its generalization ability. Our fundamental principle to split is to maximize the inclusion of representative landscape properties as possible. Specifically we visualize these 24 problems under 2D setting, and then select 12 representative problems into train set. We also provide contour map of problems in train set in Figure 8 and test set in Figure 9. Moreover, to avoid possible issue [64] coming from fixed optima which is often located in $[0,...,0]$ in current benchmark problems (this might facilitate model to overfit to this fixed point), we thus add random offset $O$ into each problems, that is to convert $y=f(x)$ into $y=f(x-O)$. This operation is inserted into both train set $\mathbb{D}_{\text{train}}$ and test set $\mathbb{D}_{\text{test}}$.
@@ A.3 Train-test split of BBOB-Noisy and Protein-Docking testsuites
  • BBOB-Noisy: this testsuits contains 30 noisy problems from COCO [57]. They are obtained by further inserting noise with different models and levels into problems in BBOB testsuits. BBOB-Noisy is characterized by its noisy nature and often used to examine robustness of certain optimizers.
  • Protein-Docking: this testsuits contains 280 instances of different protein-protein complexes [59]. These problems are characterized by rugged objective landscapes and are computationally expensive to evaluate.
@@ A.4 License of Used Open-sourced Assets
Our codebase can be accessed at https://github.com/GMC-DRL/Neur-ELA. In Table 3 we listed several open-sourced assets used in our work and their corresponding licenses.
@@
Top-level optimizer PyPop7 [65] GPL-3.0 license
MetaBBO algorithms implementation Low-level train-test workflow MetaBox [13] BSD-3-Clause license
Parallel processing Ray [55] Apache-2.0 license
ELA feature calculation pflacco [39] MIT license
@@ A.5 Control-parameters of $ES$
Fast CMAES We grid-search three key hyper-parameters in Fast CMAES, including the mean value $\mu$ and sigma value $\sigma$ of the initial Gaussian distribution used for sampling, learning rate of evolution path update $c$. We list the grid search options in Table 4 and choose the best setting according to training performance on BBOB. Besides, for other control-parameters of the Fast CMAES, we follow the default settings listed in its original paper [19].
@@
Other candidate evolution strategy variants We follows the default settings as implementations in PyPop7 [65] for other candidate top-level optimizers. We made a comparision study among SEP-CMAES [62], R1ES, RMES [63], original CMAES [18] and Fast CMAES under their default settings and finally select Fast CMAES as the default top-level optimizer of this work.

### v3

@@ 5 Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China No. 62276100, in part by the Guangdong Provincial Natural Science Foundation for Outstanding Youth Team Project No. 2024B1515040010, in part by the Guangdong Natural Science Funds for Distinguished Young Scholars No. 2022B1515020049, and in part by the TCL Young Scholars Program.
@@ A.1 Technical Details of $\Lambda_{\theta}$
Two-stage Attention Block. We construct a two-stage attention block (Ts-Attn) to aggregate optimization status information across candidate solutions and across each dimension of the decision variables. The overall computation graph of the Ts-Attn is illustrated in the right of Figure 2 in the main body, of which a basic component is the attention block ($Attn$). As illustrated in the left of Figure 2 in the main body, the $Attn$ block mainly follows designs of the original Transformer Vaswani et al. (2017), except that the layer normalization Ba et al. (2016) is used instead of batch normalization Ioffe & Szegedy (2015). Given a group of $L$ input encoding vectors $X_{in}\in\mathbb{R}^{L\times h}$, Eq. (5) details the computation of the $Attn$ block.
@@
where MHSA, LN and FF denote the multi-head self-attention Vaswani et al. (2017) (with the hidden dimension of $h$), layer normalization Ba et al. (2016) and linear feed forward layer respectively. The output $o$ holds identical shape with the input $X_{in}$. In our Ts-Attn block, we employ an $Attn$ block $Attn_{inter}$ for the first cross-solution information sharing stage, and the other $Attn$ block $Attn_{intra}$ for the second cross-dimension information sharing stage (illustrated in the right of Figure 2 in the main body. The Ts-Attn receives the input encoding $s$ of the current candidate population, and then advances the information sharing in both cross-solution and cross-dimension level. The computation is detailed in Eq. (6).
@@
At the first stage, we let the input encoding $s$ (attained from the pre-processing module) pass through $Attn_{inter}$. Since we group the encodings of the same dimension of all candidates in $s$, the $Attn_{inter}$ promotes the optimization information sharing across candidates in current population. From the first stage, we obtain a group of hidden features $H$ with the shape of $d\times m\times h$. At the second stage, we first transpose $H$ into the shape of $m\times d\times h$ to regroup all dimensions of a candidate together. We then add $cos$/$sin$ positional encoding (PE) over the transposed $H$ to inform the order of different dimensions in a candidate. We then let $H$ pass through $Attn_{intra}$ to advance the information sharing among the different dimensions within the same candidate. The output of $Attn_{intra}$ holds the shape of $m\times d\times h$. At last, we apply MeanPooling on $H_{out}$ to get the landscape feature for each candidate $F_{indiv}$ in the population, and apply a second MeanPooling on $F_{indiv}$ to get the landscape feature for the whole candidate population $F_{pop}$. We have to note that we calculate both $F_{indiv}$ and $F_{pop}$ to make our NeurELA compatible with diverse MetaBBO algorithms, which either require the landscape feature of the whole population (e.g., Wu & Wang (2022)) or require a separate landscape feature for each candidate (e.g., Sun et al. (2021)). The highly parallelizable attention-based neural-network architecture ensure the scalability of our method as the amount of the sampled candidates or the dimensions of the problem increases.
@@ A.2 Train-test split of BBOB testsuites
BBOB contains 24 synthetic problems owning various landscape properties Mersmann et al. (2011) (e.g. multi-modality, global structure, separability and etc.). Table 2 list all of the 24 problems according to their category. Due to the diversity of properties, how to split these problems into train-test set becomes a key issue to ensure the training performance and its generalization ability. Our fundamental principle to split is to maximize the inclusion of representative landscape properties as possible. Specifically we visualize these 24 problems under 2D setting, and then select 12 representative problems into train set. We also provide contour map of problems in train set in Figure 8 and test set in Figure 9. Moreover, to avoid possible issue Kudela (2022) coming from fixed optima which is often located in $[0,...,0]$ in current benchmark problems (this might facilitate model to overfit to this fixed point), we thus add random offset $O$ into each problems, that is to convert $y=f(x)$ into $y=f(x-O)$. This operation is inserted into both train set $\mathbb{D}_{\text{train}}$ and test set $\mathbb{D}_{\text{test}}$.
@@ A.3 Train-test split of BBOB-Noisy and Protein-Docking testsuites
  • BBOB-Noisy: this testsuits contains 30 noisy problems from COCO Hansen et al. (2021). They are obtained by further inserting noise with different models and levels into problems in BBOB testsuits. BBOB-Noisy is characterized by its noisy nature and often used to examine robustness of certain optimizers.
  • Protein-Docking: this testsuits contains 280 instances of different protein-protein complexes Hwang et al. (2010). These problems are characterized by rugged objective landscapes and are computationally expensive to evaluate.
@@ A.4 License of Used Open-sourced Assets
Our codebase can be accessed at https://anonymous.4open.science/r/Neur-ELA-303C. In Table 3 we listed several open-sourced assets used in our work and their corresponding licenses.
@@
Top-level optimizer PyPop7 Duan et al. (2022) GPL-3.0 license
MetaBBO algorithms implementation Low-level train-test workflow MetaBox Ma et al. (2023) BSD-3-Clause license
Parallel processing Ray Moritz et al. (2018) Apache-2.0 license
ELA feature calculation pflacco Kerschke & Trautmann (2019b) MIT license
@@ A.5 Control-parameters of $ES$
Fast CMAES We grid-search three key hyper-parameters in Fast CMAES, including the mean value $\mu$ and sigma value $\sigma$ of the initial Gaussian distribution used for sampling, learning rate of evolution path update $c$. We list the grid search options in Table 4 and choose the best setting according to training performance on BBOB. Besides, for other control-parameters of the Fast CMAES, we follow the default settings listed in its original paper Li et al. (2018).
@@
Other candidate evolution strategy variants We follows the default settings as implementations in PyPop7 Duan et al. (2022) for other candidate top-level optimizers. We made a comparision study among SEP-CMAES Ros & Hansen (2008), R1ES, RMES Li & Zhang (2017), original CMAES Hansen & Ostermeier (2001b) and Fast CMAES under their default settings and finally select Fast CMAES as the default top-level optimizer of this work.
Appendix B Additional Discussion
B.1 Training Convergence
Figure 10: Training curves of different ES baselines when training NeurELA
In NeurELA, the meta-objective as defined in Eq. (4), is non-differentiable. Hence, we train the neural network in NeurELA through neuroevolution. Such paradigm requires effective evolutionary optimizers which maintain a population of neural networks and reproduce elite offsprings iteratively according to the training objective of the neural networks. In NeurELA, we adopt Evolution Strategy (ES) since it is claimed to be more effective then other optimizers. There are many modern variants of ES method, of which we select five: Fast CMAES (Li et al., 2018), Sep-CMAES (Ros & Hansen, 2008), R1ES (Li & Zhang, 2017), RMES (Li & Zhang, 2017) and CMAES (Hansen & Ostermeier, 2001b) as candidates. We present the training curves of all five optional ES baselines under our training settings in Figure 10. The results demonstrate that the Fast CMAES we adopted for training NeurELA converges and achieves superior training effectiveness to other ES baselines.
B.2 Difference between NeurELA and Deep-ELA
Although a previous work Deep-ELA (Seiler et al., 2024) also proposed using attention-based architecture for landscape analysis, there are significant differences between our NeurELA and Deep-ELA, which we listed as below:
1. Target scenario. NeurELA is explicitly designed for MetaBBO tasks, where dynamic optimization status is critical for providing timely and accurate decision-making at the meta level. In contrast, Deep-ELA serves as a static profiling tool for global optimization problem properties and is not tailored for dynamic scenarios. NeurELA supports dynamic algorithm configuration, algorithm selection, and operator selection. In contrast, Deep-ELA’s features are restricted to static algorithm selection and configuration, limiting its adaptability in dynamic MetaBBO workflows.
2. Feature extraction workflow. First, NeurELA addresses the limited scalability of Deep-ELA for high dimensional problem. Concretely, the embedding in Deep-ELA is dependent on the problem dimension and hence the authors of Deep-ELA pre-defined a maximum dimension (50 in the original paper). To address this, NeurELA proposes a novel embedding strategy which re-organizes the sample points and their objective values to make the last dimension of the input tensor is 2 (Section 3.2). This embedding format has a significant advantage: the neural network of NeurELA is hence capable of processing any dimensional problem and any number of sample points. NeurELA enhances the information extraction through its two-stage attention-based neural network. Specifically, when processing the embedded data, Deep-ELA leverages its self-attention layers for information sharing across sample points only. In contrast, NeurELA incorporates a two-stage attention mechanism, enabling the neural network to first extract comprehensive and useful features across sample points (cross-solution attention) and then across problem dimensions (cross-dimension attention). This design helps mitigate computational bias and improve feature representation.
3. Training method. The training objective and training methodology in NeurELA and Deep-ELA are fundamentally different. Deep-ELA aims to learn a neural network that could serve as an alternative of traditional ELA. Its training objective is to minimize the contrastive loss (InfoNCE) between the outputs of its two prediction heads (termed as student head and teacher head) by gradient descent, in order to achieve invariance across different landscape augmentation on the same problem instance. In contrast, the training objective of NeurELA is to learn a neural network that could provide dynamic landscape features for various MetaBBO tasks. Specifically, its objective is to maximize the expected relative performance improvement when integrated into different MetaBBO methods. Since such relative performance improvement is not differentiable, NeurELA employs neuroevolution as its training methodology. Neuroevolution is recognized as an effective alternative to gradient descent, offering robust global optimization capabilities.
In summary, NeurELA and Deep-ELA are two totally different works with different target operating scenarios, algorithm design tasks, neural network designs and workflows, and training methodologies.
B.3 Further Interpretation Analysis
To further interpret what features have been learned by our NeurELA, we have conducted following experimental analysis to further interpret the relationship between NeurELA features and traditional ELA features, where we uses Pearson Correlation analysis to quantify the correlation between each NeurELA feature and each traditional ELA feature.Below, we explain our experimental methodology step by step:
1. We select three MetaBBO methods (LDE, RLEPSO and RL-DAS) from our training task set and employ their pre-trained models to optimize the 24 problem instances in CoCo BBOB-10D suite. Each MetaBBO method performed 10 independent runs per problem instance, with each run consisting of 500 optimization steps. Now we obtain 3*24*10 = 720 optimization trajectories, each with length 500, and the data at each step of a trajectory is the population and the corresponding objective values $\{Xs,Ys\}$.
2. Based on the obtained trajectories, we use the pre-trained NeurELA model (outputs 16 features) and the traditional ELA (we choose 32 ELA features from the traditional ELA including the Meta-model group, Convexity group, Level-Set group, Local landscape group and Distribution group) to calculate landscape features for each optimization step. After the computation, we obtain 720 landscape features time series for NeurELA and traditional ELA respectively.
[…]
Figure 11: Correlation score of NeurELA features and traditional ELA features
4. We obtain the final correlation scores of each feature pair between NeurELA and traditional ELA by averaging $r_{i,j}$ of this feature pair over the 720 time series data, $i\in\{1,2,…,16\}$, $j\in\{1,2,…32\}$. Finally we obtain a correlation matrix with 32 rows and 16 columns. We illustrate this correlation matrix by the heatmap in Figure 11. The x-axis denote 16 NeurELA features and y-axis denote 32 ELA features, a larger value denotes the two features are closely related.
From the correlation results in that Figure, we could find some relationship patterns between our NeurELA features and the traditional ELA features: a) four NeurELA features F1, F4, F8 and F16 are novel features learned by NeurELA which show weak correlation (< 0.6) with all ELA features. b) some NeurELA features show strong correlation with one particular feature group in traditional ELA, such as F3 with the Meta-model group. c) some NeurELA features show strong correlation with multiple feature groups in traditional ELA, such as F10 with Distribution group and Meta-model group. d) all NeurELA features show weak correlation with the Convexity group and Local landscape group, which might reveals these two group features are less useful for addressing MetaBBO tasks.